OpenStudy (anonymous):

Please Help... Your parents have a credit card with a balance of $3,287.90 at an interest rate of 14.5% APR. They pay $1,200.00 each month on the due date until the card is paid off. How many months does it take to pay off the card, and what is the total amount paid including interest?

OpenStudy (jskhupmang):

ok so I would make x= each month. 1200 is the slope. the interest of 3,287.90 is 476.75. So the equation would be 3,287.90=1200x+ 476.75. Now solve for x.

OpenStudy (jskhupmang):

do you know how to slove for x?

OpenStudy (anonymous):

273.99?

OpenStudy (anonymous):

@jskhupmang

OpenStudy (jskhupmang):

nope

OpenStudy (jskhupmang):

subtract 476.75 both sides

OpenStudy (jskhupmang):

then divide both sides by 1200 so that you can isolate the x.

OpenStudy (jskhupmang):

so its gonna be like this. 3,287,90-476.75=1200x+476.75-476.75 2811.15=1200x since you subtract 476.75, it is cancelled out. Then, divide both sides by 1200 to isolate the x term. 2811.15÷1200=1200x÷1200 x=2.343 So roughly it will take about 2 month and a half to payoff the balance.

OpenStudy (anonymous):

thanks bro I was clueless at least I know the formula for the other question

OpenStudy (jskhupmang):

you welcome! By the way you are the first person I have helped.

OpenStudy (anonymous):

Lol you did a great job ..

OpenStudy (jskhupmang):

thanks let me know when you need help again. I'll try my best.

OpenStudy (jskhupmang):

btw what subject is it? Finance or algebra?

OpenStudy (anonymous):

Ok and consumer math basically finance

OpenStudy (anonymous):

To calculate the monthly interest charged you divide the APR by 12 to get 1.2%. So it breaks down as follows: Month Open Bal Payment Remainder Interest Charged Closing Balance 1 3287.90 1200 2087.90 25.23 2113.13 2 2113.13 1200 913.30 11.03 924.16 3 924.16 924.16 So they pay the card of at the end of the third month having paid a total of $3.324.16 The answer does not come out formatted so take the numbers and put spaces between them until they line up under their column headings. Then it will make sense.

OpenStudy (anonymous):

did u get your answer

OpenStudy (anonymous):

hello